# Self-Supervised Millimeter Wave Indoor Localization using Tiny Neural Networks

### v1

@@ Abstract
The quasi-optical propagation of millimeter-wave (mmWave) signals enables high-accuracy localization algorithms that employ geometric approaches or machine learning models. However, most algorithms require information on the indoor environment, may entail the collection of large training datasets, or bear an infeasible computational burden for commercial off-the-shelf (COTS) devices. In this work, we propose to use tiny neural networks (NNs) to learn the relationship between angle difference-of-arrival (ADoA) measurements and locations of a receiver in an indoor environment. To relieve training data collection efforts, we resort to a self-supervised approach by bootstrapping the training of our neural network through location estimates obtained from a state-of-the-art localization algorithm. We evaluate our scheme via mmWave measurements from indoor 60-GHz double-directional channel sounding. We process the measurements to yield dominant multipath components, use the corresponding angles to compute ADoA values, and finally obtain location fixes. Results show that the tiny NN achieves sub-meter errors in 74% of the cases, thus performing as good as or even better than the state-of-the-art algorithm, with significantly lower computational complexity.
@@ I Introduction
In this work, we propose to bridge the above gap by means of a self-supervised indoor localization approach. To keep complexity minimal, we resort to a tiny NN model that exploits angle difference-of-arrival (ADoA) information as input features. Such NN models have much fewer weights compared to other architectures in the literature, thus being compatible with resource-constrained COTS hardware. To relieve the effort of collecting training data for the NN, we initially resort to a bootstrapping localization algorithm based on the same ADoA data we feed to the NN. While such algorithm will inevitably make errors when computing location estimates, it will still yield approximate location labels for ADoA readings, which can work as training data for the tiny NN. In any event, the advantage of not having to explicitly collect training data largely offsets the disadvantage of imperfect location labels in any practical scenario. As each mmWave device localizes itself, our scheme is device-centric, hence more scalable than AP-centric approaches [5].
@@
b) We reduce the training dataset collection overhead via a self-supervised approach: we train the NN model using error-prone location estimates from a bootstrapping geometric scheme that also exploits ADoAs; once trained, we only use the NN for localization.
@@ II-B Machine learning-based algorithms
Several works employed deep learning to localize mmWave devices indoors. For example, a multi-layer perceptron regression model maps the SNR fingerprints to the coordinates of a device in one dimension in [27]. Pajovic et al. used RSSI and beam index fingerprints to design probabilistic localization models [28]. and used spatial beam SNRs in [29] to learn a multi-task model for position and orientation classification. The authors in [30] and [31] proposed ResNet-inspired models for LoS and NLoS scenarios. To tackle NLoS propagation, spatial beam SNR values were used in [30], and multi-channel beam covariance matrix images in [31]. In [32], the authors used the K-nearest neighbours (K-NN) algorithm to localize unmanned aerial vehicles (UAVs) via indoor RSSI-based fingerprinting, whereas AoAs fingerprints together with RSSI measurements at reference locations were used in [33].
Although deep learning schemes yield accurate location estimates, they (a) are not easily scalable, as radio fingerprints change with the number of APs as well as the setup of the indoor environment (furniture, presence of people, etc.); (b) require knowledge of the indoor environment; (c) require large training datasets; (d) rely on models with hundreds of thousands of parameters, which may be infeasible for resource-constrained devices.
@@ III Proposed localization scheme
III-A Key idea: self-supervised tiny neural networks
Our key proposition is to localize mmWave devices via a tiny NN, trained to reproduce the behavior of an otherwise complex localization algorithm. Besides the lower computational burden that results, using a bootstrapping algorithm allows us to train the NN in a fully self-supervised fashion, with no need for the typically heavy process of collecting training data.
Fig. 1 shows the workflow of our proposed localization scheme. We start by computing ADoAs from the set of MPCs of the mmWave channel at a given location, by taking the AoA of one MPC as a reference. A tiny NN model then learns a non-linear regression function, mapping the input ADoA values to the 2D coordinates of the receiver. As training is self-supervised via a bootstrapping localization algorithm, the association between ADoA and the location where these ADoAs are observed will be inherently error-prone. Note that classification models, which would return the best-matching location from a training database for each set of measured ADoAs, would be inherently less robust to error-prone training data than regression NN models [13], besides requiring a much larger NN output stage with inherently coarse location resolution. In fact, self-supervised training with error-prone location labels is advantageous in our case, both from a practical standpoint (because it removes the need for training data collection) and from an accuracy standpoint.
@@ III-B Input features of the NN
We employ ADoA values computed from the AoAs of the MPCs of the channel extracted by the receiver, to localize mmWave devices. This removes the compass bias at the receiver, thus making localization problem invariant to receiver (RX) orientation [17]. Note that this procedure is amenable both to lab-grade devices [12, 24] and COTS devices [25, 7].
The set of MPCs that reach the receiver typically includes one LoS path and multiple NLoS paths. NLoS paths can be modeled as emanating from a (virtual) source that corresponds to the mirror image of the physical transmitter (TX) with respect to a reflective surface. We refer to these sources as virtual TXs, and collectively call physical and virtual TXs anchors. Fig. 2 illustrates one virtual anchor (VA) for the TX, representing the source of the NLoS path.
@@ III-C Clustering of the multi-path components
To compute ADoAs, we first choose the reference MPC $\bar{\bf c}_{k^{\star}}$ as the one with the least delay, where $k^{\star}=\operatorname*{arg\,min}_{k}\tau_{k}$ for $k=1,2,\cdots,|\mathcal{C}|$. We then remove $\bar{\bf c}_{k^{\star}}$ from $\mathcal{C}$, and subtract $\theta_{k^{\star}}$ from the AoAs of all remaining centroids. Finally, we wrap the ADoAs in $[0^{\circ},360^{\circ})$, and sorting the centroid set in order of increasing ADoAs. In practice, this corresponds to sweeping the azimuthal angle domain in order of increasing AoA difference from the reference MPC. WE finally form the input feature vector for the NN, by expressing all ADoAs in radians wrapped in $[\pi,\pi)$.
@@ 1 Function Rec_DBSCAN ($\bf{X}$, $\epsilon$, $\gamma$, $\eta$):
2       $\mathcal{X}\leftarrow$ DBSCAN $({\bf X},\epsilon,\gamma)$ $\mathcal{C}\leftarrow$ Centroid$(\mathcal{X})$ while $\exists\,{\bf X}_{c}\in\mathcal{X}$ of centroid ${\bf c}$ s.t. Cost$({\bf{X}}_{c},{\bf c})>\epsilon$ do
3             $\epsilon^{\prime}\leftarrow\eta\epsilon$  ,     $\gamma^{\prime}\leftarrow\eta\gamma$ foreach ${\bf X}_{c}\in\mathcal{X}$ of centroid ${\bf c}$ s.t. Cost$({\bf X}_{c},{\bf c})>\epsilon^{\prime}$  do
4                   $\mathcal{C}\leftarrow\mathcal{C}\smallsetminus\{{\bf c}\}$ $\mathcal{X}_{n}\leftarrow$ DBSCAN $({\bf X}_{c},\epsilon^{\prime},\gamma^{\prime})$ foreach  ${\bf X}_{n}\in\mathcal{X}_{n}$  do
5                         ${\bf c}_{n}\leftarrow$ Centroid$({\bf X}_{n})$ $\mathcal{C}\leftarrow$ Append$({\bf c}_{n})$
6
@@ III-D Architecture of the NN
where
$\mathcal{F}({\Theta}_{\bf x})=\mathcal{A}\big{(}{\bf W}_{4}^{T}\mathcal{A}\big{(}{\bf W}_{3}^{T}\mathcal{A}\big{(}{\bf W}_{2}^{T}\mathcal{A}({\bf W}_{1}^{T}{\Theta}_{\bf x}+{\bf b}_{1})+{\bf b}_{2}\big{)}+{\bf b}_{3}\big{)}+{\bf b}_{4}\big{)}$ (3)
is the non-linear regression function the NN applies to the ADoAs. As the number of neurons in our NNs is small, computing ${\cal F}(\cdot)$ only requires a limited number of small-size matrix multiplications and vector summations. Such operations are feasible even for computationally-constrained devices. We choose the rectified linear unit activation function (ReLU) for $\mathcal{A}(\cdot)$, and train the network using the Adam optimizer. We frame the regression as a mean-square error (MSE)minimization problem, and utilize the MSE loss function
@@
where $(x_{1},x_{2})$ are the self-supervised location labels for ADoA features from the bootstrapping algorithm, and $(\hat{x}_{1},\hat{x}_{2})$ are the location estimates yielded by the NN.
@@ III-E Hyperparameter tuning
 Node factor ($\kappa$) $\{0.6,~{}0.7,~{}0.8,~{}0.9\}$
 Dropout rate ($p$) $\{0\%,~{}10\%,~{}20\%\}$
@@
In order to optimize the NN to minimize the loss function, we automatically tune the following hyperparameters of the model: i) the node factor $\kappa$, which defines the number of neurons in the hidden layers; ii) the dropout rate $p$, which helps prevent overfitting by ignoring $p$of the NN links during the training phase; iii) the learning rate $r$; and iv) the training data batch size $b$. Table II summarizes the hyperparameter search ranges. The learning rate values range from $10^{-4}$ to $10^{-2}$ in a logarithmic progression, with ten values per decade.
@@ III-F Choice of the bootstrapping localization algorithm
In our self-supervised approach, the bootstrapping algorithm has two objectives: ($i$) to localize a device while the NN is not yet trained; ($ii$) to provide location labels that we can associate to ADoA measurements in order to train the NN. Ideally, the algorithm should work with the same input data as the NN, so that the devices does not need to measure/process additional metrics.
@@ IV-A Channel sounder
Fig. 3a displays the TX and the RX of the NIST 60 GHz 3D double-directional switched-array channel sounder. The RX embeds a circular array of 16 horn antennas having a 3D Gaussian radiation pattern with ${}^{\circ}$ beamwidth and 18.1 dBi gain. To avoid “blind spots”, the angular spacing between adjacent horns along the array’s azimuthal plane is matched to the beamwidth, i.e., subsequent horns are oriented ${}^{\circ}$ away from each other. Moreover every second horn in the circular array is oriented towards an elevation angle of ${}^{\circ}$. The resulting synthesized azimuth field-of-view of the array is ${}^{\circ}$, whereas the elevation field-of-view (FoV) is ${}^{\circ}$.
The TX is almost identical, except that it embeds a semi-circular array with 8 horns, limiting the azimuthal FoV to ${}^{\circ}$ while maintaining the same ${}^{\circ}$-elevation FoV. We synchronously trigger the TX and the RX via rubidium clocks at each end, which are also used to discipline the local oscillators. For further system details, we refer the interested readers to [40, 14].
@@ IV-C Dataset preprocessing
Figure 4: MPCs clustering using our proposed recursive approach for a sample RX location in area 4. Note that the triangles represent the cluster centroids from the initial DBSCAN procedure and the stars represent the centroids obtained after recursive reclustering. (a) Reference ray traces of the LoS and first-order MPCs from TX-1 to the RX. (b) Clustering of the MPCs from TX-1. (c) Clustering of the MPCs from all four TXs.
@@ V-A Illustration of the MPCs clustering operation
In Fig. 4b, we mark cluster centroids as black triangles (after the initial DBSCAN pass) and stars (after the recursive clustering procedure). As a first result, we observe that hierarchical reclustering is needed otherwise, e.g., the earliest arrivals having an AoA of $300^{\circ}$ and above (in dark and light blue towards the top of the picture) would be mistaken as a single arrival. Each centroid can be interpreted as a dominant MPC. We numbered the most significant ones to illustrate that identified clustering include all key MPCs expected from the ray trace in panel LABEL:sub@fig:rayTrace. The centroid corresponding to the smallest MPC delay represents the LoS path (ray 1 in Fig. 4a), as confirmed by its AoA of about $250^{\circ}$. Similarly, MPC 2 has slightly larger AoA and delay features with respect to MPC 1; conversely, MPC 4 has a comparatively longer path length and an AoA of about $100^{\circ}$. The rest of the MPCs from TX-1 are annotated based on the scheme proposed in Section III-C.
From the azimuthal AoA features of the cluster centroids, we form the set of ADoAs values for the localization algorithm. Fig. 4c shows the centroids resulting from applying the clustering process to MPCs obtained from all of the four available transmitters.

### v2

# Algorithm-Supervised Millimeter Wave Indoor Localization using Tiny Neural Networks

@@ Abstract
The quasi-optical propagation of millimeter-wave (mmWave) signals enables high-accuracy localization algorithms that employ geometric approaches or machine learning models. However, most algorithms require information on the indoor environment, may entail the collection of large training datasets, or bear an infeasible computational burden for commercial off-the-shelf (COTS) devices. In this work, we propose to use tiny neural networks (NNs) to learn the relationship between angle difference-of-arrival (ADoA) measurements and locations of a receiver in an indoor environment. To relieve training data collection efforts, we resort to an algorithm-supervised approach by bootstrapping the training of our neural network through location estimates obtained from a state-of-the-art localization algorithm. We evaluate our scheme via mmWave measurements from indoor 60-GHz double-directional channel sounding. We process the measurements to yield dominant multipath components, use the corresponding angles to compute ADoA values, and finally obtain location fixes. Results show that the tiny NN achieves sub-meter errors in 74% of the cases, thus performing as good as or even better than the state-of-the-art algorithm, with significantly lower computational complexity.
@@ I Introduction
In this work, we propose to bridge the above gap by means of an algorithm-supervised indoor localization approach. To keep complexity minimal, we resort to a tiny NN model that exploits angle difference-of-arrival (ADoA) information as input features. Such NN models have much fewer weights compared to other architectures in the literature, thus being compatible with resource-constrained COTS hardware. To relieve the effort of collecting training data for the NN, we initially resort to a bootstrapping localization algorithm based on the same ADoA data we feed to the NN. While such algorithm will inevitably make errors when computing location estimates, it will still yield approximate location labels for ADoA readings, which can work as training data for the tiny NN. In any event, the advantage of not having to explicitly collect training data largely offsets the disadvantage of imperfect location labels in any practical scenario. As each mmWave device localizes itself, our scheme is device-centric, hence more scalable than AP-centric approaches [5].
@@
b) We reduce the training dataset collection overhead via an algorithm-supervised approach: we train the NN model using error-prone location estimates from a bootstrapping geometric scheme that also exploits ADoAs; once trained, we only use the NN for localization.
@@
d) We complement our evaluation with additional simulation results from a complex-shaped room environment.
@@ II-B Machine learning-based algorithms
Several works employed deep learning to localize mmWave devices indoors. For example, a multi-layer perceptron regression model maps the SNR fingerprints to the coordinates of a device in one dimension in [27]. Pajovic et al. used RSSI and beam index fingerprints to design probabilistic localization models [28] and used spatial beam SNRs in [29] to learn a multi-task model for position and orientation classification. The authors in [30] and [31] proposed ResNet-inspired models for LoS and NLoS scenarios. To tackle NLoS propagation, spatial beam SNR values were used in [30], and multi-channel beam covariance matrix images in [31]. In [32], the authors used the K-nearest neighbours (K-NN) algorithm to localize unmanned aerial vehicles (UAVs) via indoor RSSI-based fingerprinting, whereas AoAs fingerprints together with RSSI measurements at reference locations were used in [33].
Although deep learning schemes yield accurate location estimates, they (a) are not easily scalable, as radio fingerprints change with the number of APs as well as the setup of the indoor environment (furniture, presence of people, etc.); (b) require knowledge of the indoor environment; (c) require large training datasets; (d) require fusing multiple mmWave signal parameters to form rich input feature vectors for the models; (e) rely on models with hundreds of thousands of parameters, which may be infeasible for resource-constrained devices.
@@ III Proposed localization scheme
III-A Key idea: algorithm-supervised tiny neural networks
Our key proposition is to localize mmWave devices via a tiny NN, trained to reproduce the behavior of an otherwise complex localization algorithm. Besides the lower computational burden that results, using a bootstrapping algorithm allows us to train the NN in a fully algorithm-supervised fashion, with no need for the typically heavy process of collecting training data.
Fig. 1 shows the workflow of our proposed localization scheme. We assume that the SNR between the transmitter and receiver is sufficient to achieve over-the-air synchronization. We start by computing ADoAs from the set of MPCs of the mmWave channel at a given location, by taking the AoA of one MPC as a reference. A tiny NN model then learns a non-linear regression function, mapping the input ADoA values to the 2D coordinates of the receiver. As training is algorithm-supervised via a bootstrapping localization algorithm, the association between ADoA and the location where these ADoAs are observed will be inherently error-prone. Note that classification models, which would return the best-matching location from a training database for each set of measured ADoAs, would be inherently less robust to error-prone training data than regression NN models [13], besides requiring a much larger NN output stage with inherently coarse location resolution. In fact, algorithm-supervised training with error-prone location labels is advantageous in our case, both from a practical standpoint (because it removes the need for training data collection) and from an accuracy standpoint.
@@ III-B Input features of the NN
We employ ADoA values computed from the AoAs of the MPCs of the channel extracted by the receiver, to localize mmWave devices.333 We neglect the MPC delay as the input to both the schemes, as most commercially deployed indoor mmWave systems such as IEEE 802.11ad/ay wireless LAN (WLAN) systems require additional CSI pre-processing steps in order to retrieve the MPC delays, unlike the AoAs. While there exist standards for estimating the time of flight, e.g., via the fine time measurement (FTM) protocol (IEEE 802.11-REVmc), these methods generally estimate the delay corresponding to the strongest path between the transmitter (TX) and the receiver (RX), which can be either a LoS or NLoS path. This removes the compass bias at the receiver, thus making localization problem invariant to RX orientation [17]. Note that this procedure is amenable both to lab-grade devices [12, 24] and COTS devices [25, 7].
The set of MPCs that reach the receiver typically includes one LoS path and multiple NLoS paths. NLoS paths can be modeled as emanating from a (virtual) source that corresponds to the mirror image of the physical TX with respect to a reflective surface. We refer to these sources as virtual TXs, and collectively call physical and virtual TXs anchors. Fig. 2 illustrates one virtual anchor (VA) for the TX, representing the source of the NLoS path.
@@ III-C Clustering of the multi-path components
We remark that the average complexity of a regular DBSCAN procedure is $\mathcal{O}(N\log N)$, where $N$ is the total number of MPCs obtained from each TX (in the order of few hundreds). On the contrary, the complexity of recursive DBSCAN procedure is $\mathcal{O}(N\log N+K\cdot m\log m)$, where $K$ is the upper-bound average number of clusters to be re-clustered and $m$ is the average number of MPCs in such cluster. As our proposed procedure only breaks down those clusters that have a total associated cost exceeding $\epsilon$, $K$ and $m$ are significantly less than $N$. Thus, the increase in the average complexity remains relatively insignificant.
Dynamic changes in the environment can affect small-scale fading parameters such as phase and MPC delay [40]. However, recursively clustering these MPCs and subsequently computing the centroids helps average-out these fading effects. As these centroids eventually correspond to the AoAs (and thus, the ADoAs) of the dominant MPCs, our localization schemes can thus be resilient to dynamic environments.
To compute ADoAs, we first choose the reference MPC $\bar{\bf c}_{k^{\star}}$ as the one with the least delay, where $k^{\star}=\operatorname*{arg\,min}_{k}\tau_{k}$ for $k=1,2,\cdots,|\mathcal{C}|$. We then remove $\bar{\bf c}_{k^{\star}}$ from $\mathcal{C}$ and sort the AoAs of the remaining centroids in increasing order. In practice, this corresponds to sweeping the entire azimuthal angle domain in increasing order of the AoAs (as seen from the global reference frame). We then subtract $\theta_{k^{\star}}$ from the sorted AoAs. We finally form the input feature vector for the NN, by expressing all ADoAs in radians wrapped in $[-\pi,\pi)$.
1
@@ 2 Function Rec_DBSCAN ($\bf{X}$, $\epsilon$, $\gamma$, $\eta$):
3       $\mathcal{X}\leftarrow$ DBSCAN $({\bf X},\epsilon,\gamma)$
4       $\mathcal{C}\leftarrow$ Centroid$(\mathcal{X})$
5
6      while $\exists\,{\bf X}_{c}\in\mathcal{X}$ of centroid ${\bf c}$ s.t. Cost$({\bf{X}}_{c},{\bf c})>\epsilon$ do
7             $\epsilon^{\prime}\leftarrow\eta\epsilon$  ,     $\gamma^{\prime}\leftarrow\eta\gamma$
8             foreach ${\bf X}_{c}\in\mathcal{X}$ of centroid ${\bf c}$ s.t. Cost$({\bf X}_{c},{\bf c})>\epsilon^{\prime}$  do
9                   $\mathcal{C}\leftarrow\mathcal{C}\smallsetminus\{{\bf c}\}$
10                   $\mathcal{X}_{n}\leftarrow$ DBSCAN $({\bf X}_{c},\epsilon^{\prime},\gamma^{\prime})$
11                   foreach  ${\bf X}_{n}\in\mathcal{X}_{n}$  do
12                         ${\bf c}_{n}\leftarrow$ Centroid$({\bf X}_{n})$
13                         $\mathcal{C}\leftarrow$ Append$({\bf c}_{n})$
14
15
@@ III-D Architecture of the NN
where {dmath} F(Θ_x) = A(W_4^TA( W_3^T A( W_2^T A(W_1^TΘ_x+b_1) + b_2) +b_3 ) +b_4 ) is the non-linear regression function the NN applies to the ADoAs. As the number of neurons in our NNs is small, computing ${\cal F}(\cdot)$ only requires a limited number of small-size matrix multiplications and vector summations. Such operations are feasible even for computationally-constrained devices. We choose the rectified linear unit activation function (ReLU) for $\mathcal{A}(\cdot)$, and train the network using the Adam optimizer. We frame the regression as a mean-square error (MSE)minimization problem, and utilize the MSE loss function
@@
where $(x_{1},x_{2})$ are the algorithm-supervised location labels for ADoA features from the bootstrapping algorithm, and $(\hat{x}_{1},\hat{x}_{2})$ are the location estimates yielded by the NN.
@@ III-E Hyperparameter tuning
Node factor ($\kappa$) $\{0.6,~{}0.7,~{}0.8,~{}0.9,~{}1\}$
Dropout rate ($p$) $\{0\%,~{}10\%,~{}15\%,~{}20\%\}$
@@
In order to optimize the NN to minimize the loss function, we automatically tune the following hyperparameters of the model: i) the node factor $\kappa$, which defines the number of neurons in the hidden layers; ii) the dropout rate $p$, which helps prevent overfitting by ignoring $p$of the NN links during the training phase; iii) the learning rate $r$; and iv) the training data batch size $b$. Table II summarizes the hyperparameter search ranges. The learning rate values range from $10^{-4}$ to $10^{-2}$ in a logarithmic progression, with ten values per decade. The optimal hyperparameters for the NN model are chosen from given set after an exhaustive grid search technique.
@@ III-F Choice of the bootstrapping localization algorithm
In our algorithm-supervised approach, the bootstrapping algorithm has two objectives: ($i$) to localize a device while the NN is not yet trained; ($ii$) to provide location labels that we can associate to ADoA measurements in order to train the NN. Ideally, the algorithm should work with the same input data as the NN, so that the devices does not need to measure/process additional metrics.
@@ IV-A Channel sounder
Fig. 3a displays the TX and the RX of the NIST 60 GHz 3D double-directional switched-array channel sounder. The RX embeds a circular array of 16 horn antennas having a 3D Gaussian radiation pattern with 22.5 beamwidth and 18.1 dBi gain. To avoid “blind spots”, the angular spacing between adjacent horns along the array’s azimuthal plane is matched to the beamwidth, i.e., subsequent horns are oriented 22.5 away from each other. Moreover every second horn in the circular array is oriented towards an elevation angle of 22.5. The resulting synthesized azimuth field-of-view of the array is 360, whereas the elevation field-of-view (FoV) is 45.
The TX is almost identical, except that it embeds a semi-circular array with 8 horns, limiting the azimuthal FoV to 180 while maintaining the same 45-elevation FoV. We synchronously trigger the TX and the RX via rubidium clocks at each end, which are also used to discipline the local oscillators. For further system details, we refer the interested readers to [41, 14].
@@ IV-C Dataset preprocessing
Figure 4: MPCs clustering using our proposed recursive approach for a sample RX location in area 4. Note that the triangles represent the cluster centroids from the initial DBSCAN procedure and the squares represent the centroids obtained after recursive clustering. (a) Reference ray traces of the LoS and first-order MPCs from TX-1 to the RX. (b) Clustering of the MPCs from TX-1.
@@ V-A Illustration of the MPCs clustering operation
In Fig. 4b, we mark cluster centroids as black triangles (after the initial DBSCAN pass) and squares (after the recursive clustering procedure). As a first result, we observe that hierarchical reclustering is needed otherwise, e.g., the earliest arrivals having an AoA of $300^{\circ}$ and above (in dark and light blue towards the top of the picture) would be mistaken as a single arrival. Each centroid can be interpreted as a dominant MPC. We numbered the most significant ones to illustrate that identified clustering include all key MPCs expected from the ray trace in panel LABEL:sub@fig:rayTrace. The centroid corresponding to the smallest MPC delay represents the LoS path (ray 1 in Fig. 4a), as confirmed by its AoA of about $250^{\circ}$. Similarly, MPC 2 has slightly larger AoA and delay features with respect to MPC 1; conversely, MPC 4 has a comparatively longer path length and an AoA of about $100^{\circ}$. The rest of the MPCs from TX-1 are annotated based on the scheme proposed in Section III-C.
From the azimuthal AoA features of the cluster centroids, we form the set of ADoAs values for the localization algorithm.